Calculate the force between charges 4x10^-8 C and 1.8x10^-6 C if they are 3.5 m apart


1
Expert's answer
2021-05-19T17:03:04-0400

The force between these charges can be calculated by Coulomb's law:


"F=\frac{1}{4\pi\epsilon_0}\frac{q_1q_2}{R^2},\\\space\\
F=\frac{1}{4(3.14)(8.85·10^{-12})}\frac{(4·10^{-8})·(1.8·10^{-6})}{3.5^2}=5.28·10^{-5}\text{ N}."